# Improving Precision in Kinematic Weak Lensing with MIRoRS: Model-Independent Restoration of Reflection Symmetries

### v1

@@ 3.1 Overview
3.2 Direct Shear Mapping
The Direct Shear Mapping (DSM) technique was first developed by dBD15. By exploiting the effect of weak lensing on the intrinsic symmetry of the velocity field of a galaxy, they were able to fit for the shear parameter, $\gamma_{\times}$, in a completely model-independent fashion. While conceptually similar, our implementation is materially different, with an emphasis on accurately capturing the model uncertainties.
The velocity field of a rotationally-supported disk galaxy is expected to have intrinsic symmetry across its major axis and antisymmetry across its minor, with the major and minor axes being orthogonal. After lensing, the axes are no longer orthogonal, but have a separation given by the difference between Equation 18 and Equation 19,
$\theta_{\perp}\approx\frac{\pi}{2}\pm 2\gamma_{\times}.$ (24)
Fitting for $\gamma_{\times}$ with only this separation of the axes is possible, but does not fully utilize the information available. Instead, DSM uses velocity data across the entire face of the galaxy to quantify the loss of symmetry.
Figure 3: Illustration of the nonparametric fitting algorithm. Starting in the detector frame (light background), data (green) is first masked around a central pixel. To create a model, the raw data is translated to the detector origin and inverse shear is applied, bringing the image into the source plane (dark background). A clockwise rotation orients the galaxy axes with the coordinate axes. Two symmetry transformations are done independently; a reflection about the major axis (red) and an inversion about the minor (blue) using Equation 33. Rotation and shear are then reapplied to bring the images back to the source plane where they are translated back to the original position. The two separate images are combined into a single model (purple) and masked with the same mask as applied to the data. The difference between the model and the data produces the residual (orange). Here, the data has an applied shear, $\gamma_{\times}=0.15$ and position angle $\phi=30^{\circ}$ while the model transformations use $\gamma_{\times}=0.12$ and $\phi=29^{\circ}$ so a non-zero residual is produced.
@@ 3.2 Direct Shear Mapping
    The field is first translated from its observed position, $(x_{c},y_{c})$, to the detector origin. Future operations (rotation, shear) are implemented about this origin. Note that $(x_{c},y_{c})$ are then fitting parameters, and not simply an observation.
@@
    Next, we apply a shear transformation to restore the unlensed fields in the source coordinates, $[x,y]$. With the correct parameters for $(x_{c},y_{c})$ and $\gamma_{\times}$, the intrinsic symmetry of the field will be restored at this point.
@@
    In order to properly implement the two symmetry transformations, the galaxy axes must be aligned with the coordinate axes. Thus, the unsheared fields are rotated by a parameter for the position angle in the source plane, $\phi$.
@@
    Two independent symmetry transformations are applied to the velocity field: a reflection across the major axis and an inversion across the minor axis. Each transformation is performed separately, resulting in distinct model fields for the major and minor axes. To account for the antisymmetry across the minor axis, the inversion consists of a reflection followed by a sign change. To derive the expression for the minor-axis model field, the central velocity parameter, $V_{c}$, is subtracted from both the reflected and model fields, and the sign is inverted, giving $V_{\text{model}}-V_{c}=-(V_{\text{reflect}}-V_{c})$, which simplifies to $V_{\text{model}}=2V_{c}-V_{\text{reflect}}$.
@@
    These transformed fields should be identical to the original field if the correct parameters are used in the transformations (assuming, of course, intrinsic symmetry). The fields are averaged into a single model. This model is then transformed back into the detector plane by applying the inverse of each operation, rotation, shear, translation.
@@
3.3 Photometric Axes
To constrain the shear parameter from photometric data, we need to calculate the difference between the detector-plane kinematic and photometric axes. The photometric position angle in the detector plane is directly observable. Fitting for the kinematic position angle, on the other hand, is more troublesome. These complications are dealt with in Krajnovic et al. (2006), however, uncertainties in kinematic position angle from this method range from $3^{\circ}$ for rotational velocities exceeding $100$ km/s to $17^{\circ}$ for rotational velocities lower than $100$ km/s (Krajnović et al., 2011).
@@ 3.3 Photometric Axes
Combining with Equation 20 and solving for the photometric major axis position angle, we have
@@ 3.4 Optimizer
We also implemented MCMC analysis with emcee (Foreman-Mackey et al., 2013) and found similar results, but at increased computation time. Because we are interested in ensemble variances rather than the variances within the fit of a particular halo, we opted for the computational efficiency of dual_annealing. MCMC did help to provide valuable information about the nature of the likelihood surfaces we were exploring and would be a useful tool in analyzing individual halos and their associated uncertainties.
3.5 Fitting Algorithm
3.5.1 Kinematic Fitting
The function to be optimized is the negative log-likelihood. This likelihood is calculated from both the kinematic and photometric data with their associated uncertainties. For the kinematic calculation, we start with the observed velocity field and uncertainty map. The data is pre-processed to ensure all galaxies in the analysis have the same rotation direction and any missing pixels have their uncertainty set to infinity.
The likelihood depends on the residual between a transformed model and the data so, first, a data region must be defined. The transformations to create the model, particularly the shear transformation, distort the field in such a manner that pixels with no data from outside the field can be brought into the model. This necessitates using a mask region small enough that reasonable transformation parameters cannot introduce these NaN pixels. This region is selected by taking the pixels within a prescribed radius of the center of the field. Note, while the centroid location is a fitting parameter, the data region must be established prior to the fitting so that all residual calculations done by the optimizer take the difference from the same set of data.
@@ 3.5.1 Kinematic Fitting
These matrices are combined into two overall transformation operations and applied to both the data and the error maps. Applying the transformations in a single step requires only a single interpolation, rather than an interpolation for each operation. These operations do not generally commute, so care must be taken to perform the operations in the correct sequence. First, the maps are translated to the detector origin and inverse shear is applied to bring the observed field back to the source plane. Next, a clockwise rotation to align the major axis with the x-axis. The fields are then reflected across the coordinate axes, creating two sets of velocity and error maps. Each of these maps is then rotated counterclockwise and the shear re-applied. This brings the images back to the detector plane where they are translated back to the original centroid and averaged into a single model to compute a residual. Combined, these transformations are
@@
$\mathcal{M}_{\text{minor}}$ $=\mathcal{A}\mathcal{R}\mathcal{T_{\text{y}}}\mathcal{R}^{-1}\mathcal{A}^{-1}.$ (31)
@@
$\sigma_{\text{model}}=\sqrt{\sigma_{\text{x}}^{2}+\sigma_{\text{y}}^{2}}.$ (35)
The overall per-pixel uncertainty is the quadrature sum of uncertainties in both model and data,
$\sigma_{i}=\sqrt{\sigma_{\text{model}}^{2}+\sigma_{\text{data}}^{2}}.$ (36)
The fields are then masked at the same position as the data and a per-pixel $\chi^{2}$ is computed,
$\chi^{2}_{i}=\frac{\left(V_{i\text{,model}}-V_{i\text{,data}}\right)^{2}}{\sigma_{i}^{2}}.$ (37)
With a per-pixel $\chi^{2}$, our likelihood function to be maximized is
$\mathcal{L}=\prod_{i}\frac{1}{\sqrt{2\pi\sigma_{i}^{2}}}\mathrm{e}^{-\chi_{i}^{2}/2}.$ (38)
Taking the log-likelihood gives
@@
This differs from dBD15 in that no $180^{\circ}$ rotation operation is done and the $\chi^{2}$ is computed from an averaged model instead of simply summing over total pixels. On the first issue of the rotation, we can express such a rotation by
@@
This is simply the negative of the identity matrix, the full transformation matrix is then
@@
This holds for any shear parameter or position angle, thus rotations are independent of shear. Rotation operations could help constrain the other parameters in the problem—the center position and central velocity—but this would need to be done separately from the fitting of the shear parameter and implemented as a prior. Including the rotation operation in our model would increase the noise without adding any information.
On the second issue of creating an averaged model versus separately computing a $\chi^{2}$ for the individual fields produced by each reflection, we find that summing over the pixels of the fields separately underestimates the error due to the individual fields not being truly independent. Consider the case of the reflection across the major axis alone. The residual is the difference between data and model, but the model itself is built from the data. If we look at a simplified, 2-pixel image, with one upper and one lower pixel, identified with subscripts U and L respectively, the $\chi^{2}$ value (using unit uncertainty in the data) would be
@@
Alternatively, rather than reflecting the data, the operation can be thought of as a folding of the data across its axis demonstrating that the number of true comparisons is only half the pixels. The same concept holds true for the reflection across the minor axis. By summing over the entirety of pixels in each transformed field separately, the number of independent data points is artificially inflated. Instead, we choose to average over two transformations which creates a unique comparison for each pixel in the image.
3.5.2 Photometric Fitting
@@
where $M_{ij}=\sum_{x,y}x^{i}y^{j}I(x,y)$ denotes the raw moment and $I(x,y)$ is the intensity at pixel coordinates $(x,y)$.
@@
With the axis ratio, we can calculate a model value for the photometric position angle from Equation 26. The axis ratio measurement introduces uncertainty in the model value which must be accounted for. The uncertainty in the model term is calculated through standard propagation of errors
@@
Combining this term with the uncertainty in the observed position angle gives the total error in the photometric angle,
$\sigma_{\theta^{\prime}}^{2}=\sigma_{\theta^{\prime}\text{,model}}^{2}+\sigma_{\theta^{\prime}\text{,obs}}^{2}.$ (49)
The uncertainties in observed position angle and axis ratio are determined directly from the distribution of position angles and axis ratios in our data sample, as shown in Subsection 4.1. A $\chi^{2}$ value is then calculated
@@
3.6 Data Selection
3.6.1 Idealized Mock Data
To characterize the performance of our nonparametric approach, we first create a set of idealized mock data for both the velocity and intensity fields. The intensity field is described by an exponentially decaying profile,
@@ 3.6.1 Idealized Mock Data
For the velocity field we model an infinitely thin disk with an arctangent profile with the major axis aligned along the $x$-axis,
@@
where $r_{0}$ is the scale radius independent of the intensity, $V_{\text{max}}$ is the asymptotic velocity, and $V_{c}$ the central velocity of the galaxy.
@@
Both fields are then interpolated from the source plane coordinates onto a grid which has been transformed by a shear parameter, $\gamma_{\times}$, major axis position angle, $\phi$, and inclination $i$.
3.6.2 Illustris TNG
[…]
The subhalos were all oriented according to the principal axis of the moment of inertia tensor calculated from the stellar particles. The longest principal axis was aligned along the coordinate x-axis as the major axis while the second largest principal axis was aligned as the minor axis. The galaxies were also arranged to have a clockwise rotation.
3.7 Outlier Rejection
For a well-fit model, residuals are expected to be primarily due to uncorrelated noise. Spatially correlated patterns in residuals can be the result of a poorly fit parameter or from unmodeled physical effects such as spiral arms, bars, or warps. We use a Global Moran’s I-test for spatial autocorrelation (Moran, 1948) to quantify the extent to which these patterns are present in the residuals, identifying halos with a poor fit. Examining residuals for spatial autocorrelation is not a new technique (Goodchild, 1986), but has been largely limited to fields outside of astronomy. In using the Moran’s I-test, we seek to quantify the effectiveness of our modeling, but also to explore the utility of spatial autocorrelation in characterizing residuals.
To implement the Moran’s I-test, we first compute the residuals by subtracting the model values from the observed values. These residuals are then normalized by their combined errors, resulting in a standardized residuals image. The Moran’s I statistic is then calculated to measure spatial autocorrelation, using the package PySAL.esda (Rey & Anselin, 2007). Specifically, we use the Queen contiguity method (named for the directions a Queen may move in chess) to create a spatial weight matrix, where each element $w_{ij}$ is 1 if locations $i$ and $j$ are neighbors (sharing a border or vertex) and 0 otherwise. The Moran’s I statistic is given by

### v2

@@ III.1 Overview
III.2 Direct Shear Mapping
The MIRoRS method is based largely upon the direct shear mapping (DSM) technique first developed in dBD15. By exploiting the effect of weak lensing on the intrinsic symmetry of the velocity field of a galaxy, they were able to fit for the shear parameter, $\gamma_{\times}$, in a completely model-independent fashion. While conceptually similar, our implementation is materially different, with an emphasis on accurately capturing the model uncertainties.
@@ III.2 Direct Shear Mapping
The velocity field of a rotationally supported disk galaxy is expected to have intrinsic symmetry across its major axis and antisymmetry across its minor, with the major and minor axes being orthogonal. After lensing, the axes are no longer orthogonal, but have a separation given by the difference between Equation 9 and Equation 10,
$\theta_{\perp}\approx\frac{\pi}{2}\pm 2\gamma_{\times}.$ (16)
Figure 3: Illustration of the MIRoRS algorithm. Starting in the detector frame (light background), data (green) is first masked around a central pixel. To create a model, the raw data is translated to the detector origin and inverse shear is applied, bringing the image into the source plane (dark background). A clockwise rotation orients the galaxy axes with the coordinate axes. Two symmetry transformations are done independently; a reflection about the major axis (red) and an inversion about the minor (blue) using Equation 24. Rotation and shear are then reapplied to bring the images back to the source plane where they are translated back to the original position. The two separate images are combined into a single model (purple) and masked with the same mask as applied to the data. The difference between the model and the data produces the residual (orange). Here, the data has an applied shear, $\gamma_{\times}=0.15$ and position angle $\phi=30^{\circ}$ while the model transformations use $\gamma_{\times}=0.12$ and $\phi=29^{\circ}$ so a nonzero residual is produced. Note, the parameter values and mask sizing were chosen to provide qualitative insight and are not necessarily indicative of typical values.
@@
    The field is first translated from its observed position to the detector origin by fitting parameters $x_{c}$ and $y_{c}$. Future operations (rotation, shear) are implemented about this origin.
@@
    Next, we apply a shear transformation to restore the unlensed fields in the source coordinates. With the correct parameters for $(x_{c},y_{c})$ and $\gamma_{\times}$, the intrinsic symmetry of the field will be restored at this point.
@@
    To properly implement the symmetry transformations, the galaxy axes must be aligned with the coordinate axes. Thus, the unsheared fields are rotated by a parameter for the position angle in the source plane, $\phi$.
@@
    Two independent symmetry transformations are applied to the velocity field: a reflection across the major axis and an inversion across the minor axis. To account for the antisymmetry across the minor axis, the inversion consists of a reflection followed by a sign change.
@@
    These separate transformed fields should be identical to the original field if the correct parameters are used in the transformations (assuming, of course, intrinsic symmetry). The fields are averaged into a single model. This model is then transformed back into the detector plane by applying the inverse of each operation, rotation, shear, translation.
@@
III.3 Photometric Axes
To constrain the shear parameter from photometric data, we need to calculate the difference between the detector-plane kinematic and photometric axes. The photometric position angle in the detector plane is directly observable. Fitting for the kinematic position angle, on the other hand, is more troublesome. These complications are dealt with in Krajnovic et al. (2006); however, uncertainties in kinematic position angle from this method range from $3^{\circ}$ for rotational velocities exceeding $100\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$ to $17^{\circ}$ for rotational velocities lower than $100\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$(Krajnović et al., 2011).
@@ III.3 Photometric Axes
Combining with Equation 11 and solving for the photometric major-axis position angle, we have
@@ III.4 Optimizer
We also implemented Markov Chain Monte Carlo (MCMC) analysis with emcee (Foreman-Mackey et al., 2013) and found similar results, but at increased computation time. Because we are interested in ensemble variances rather than the variances within the fit of a particular halo, we opted for the computational efficiency of dual_annealing. MCMC did help to provide valuable information about the nature of the likelihood surfaces we were exploring and would be a useful tool in analyzing individual halos and their associated uncertainties.
III.5 Fitting Algorithm
III.5.1 Kinematic Fitting
The function to be optimized is the negative log-likelihood. This likelihood is calculated from both the kinematic and photometric data with their associated uncertainties. For the kinematic calculation, we start with the observed velocity field and uncertainty map. The data is pre-processed to ensure all galaxies in the analysis have the same rotation direction, and any missing pixels have their uncertainty set to infinity.
The likelihood depends on the residual between a transformed model and the data, so, first, a data region must be defined. The transformations to create the model, particularly the shear transformation, distort the field in such a manner that pixels with no data from outside the field can be brought into the model. This necessitates using a mask region small enough that reasonable transformation parameters cannot introduce these NaN pixels. This region is selected by taking the pixels within a prescribed radius of the center of the field. Note, while the centroid location is a fitting parameter, the data region must be established prior to the fitting so that all residual calculations done by the optimizer take the difference from the same set of data.
@@ III.5.1 Kinematic Fitting
These matrices are combined into two overall transformation operations and applied to both the data and the error maps. Applying the transformations in a single step requires only a single interpolation, rather than an interpolation for each operation. These operations do not generally commute, so care must be taken to perform the operations in the correct sequence. First, the maps are translated to the detector origin, and inverse shear is applied to bring the observed field back to the source plane. Next, a clockwise rotation is implemented to align the major axis with the $x$-axis. The fields are then reflected across the coordinate axes, creating two sets of velocity and error maps. Each of these maps is then rotated counterclockwise and the shear reapplied. This brings the images back to the detector plane where they are translated back to the original centroid and averaged into a single model to compute a residual. Combined, these transformations are
@@
$\mathcal{M}_{\text{minor}}$ $=\mathcal{A}\mathcal{R}\mathcal{T_{\text{y}}}\mathcal{R}^{-1}\mathcal{A}^{-1}.$
@@
$\sigma_{\text{model}}=\frac{\sqrt{\sigma_{\text{x}}^{2}+\sigma_{\text{y}}^{2}}}{2}.$ (26)
The overall per-pixel uncertainty is the quadrature sum of uncertainties in both model and data. The fields are then masked at the same position as the data, and a per-pixel $\chi^{2}$ and likelihood are computed. The log-likelihood function is then
@@
Figure 4: Fisher information density for each fitted parameter of an ideal model ($\gamma_{\times}=\phi=x_{c}=y_{c}=v_{c}=0$), computed independently for the major-axis reflection, the minor-axis reflection, and the $180^{\circ}$ rotation operations. Each panel shows the spatial distribution of $(\partial v/\partial\theta)^{2}/\sigma^{2}$ for a single parameter. The two reflection operators exhibit identical sensitivity to the shear parameter $\gamma_{\times}$, whereas the rotation operation provides no information on shear or position angle. The major-axis reflection constrains $y_{c}$ while the minor-axis reflection is sensitive to $x_{c}$ and $v_{c}$.
@@
This differs from dBD15 in that no $180^{\circ}$ rotation operation is done, and the $\chi^{2}$ is computed from an averaged model instead of simply summing over total pixels. On the first issue of the rotation, we can express such a rotation by
@@
This is simply the negative of the identity matrix; the full transformation matrix is then
@@
This holds for any shear parameter or position angle; thus, rotations are independent of shear. This is demonstrated in Figure 4, where we compute the Fisher information density for an ideal model with parameters $\gamma_{\times}=\phi=x_{c}=y_{c}=v_{c}=0$ and evaluate the response of the velocity field to perturbations in each parameter. The resulting maps demonstrate that the two reflections contain identical sensitivity to the shear parameter $\gamma_{\times}$ while rotations carry no information about the shear, position angle, or central velocity.
The Fisher maps further show how the remaining parameters enter the problem: the major-axis reflection is sensitive to displacements perpendicular to that axis, constraining $y_{c}$, whereas the minor-axis reflection constrains $x_{c}$ as well as $v_{c}$. The two reflections show more complicated sensitivities when the parameters are not zero. This is demonstrated in the correlation matrices given in Figure 5, where the top row of panels shows the correlations for the ideal model used in Figure 4 and the bottom row shows the correlations for a model created with nonzero parameters $\gamma_{\times}=0.01$, $\phi=5^{\circ}$, $x_{c}=y_{c}=1$, and $v_{c}=15\ \mathrm{km\,s^{-1}}$. In the top row, both reflections show a correlation between $\gamma_{\times}$ and $\phi$ while the rotation operation, with no constraining power on $\gamma_{\times}$ or $\phi$, does not show this coupling. The minor-axis operation also shows a coupling between $x_{c}$ and $v_{c}$, as expected. Both the Fisher density and correlation matrices suggest that the reflection operations carry the same information (apart from $v_{c}$). However, with nontrivial model parameters, it is evident that the operations have differing sensitivities and constraining power. Cross-correlations between parameters emerge differently for each operation, with the major-axis reflection remaining insensitive to $v_{c}$ and rotations only sensitive to centroid position. These distinct but complementary sensitivities motivate the use of the two reflections for parameter inference while omitting the rotation, which would contribute noise without adding information about $\gamma_{\times}$.
Figure 5: Correlation matrices for the fitted parameters $(\gamma_{\times},\ \phi,\ x_{c},\ y_{c},\ v_{c})$ derived from the Fisher–matrix covariance for the three symmetry operations. The top row corresponds to the ideal model used in Figure 4, where the expected parameter couplings appear cleanly: both reflections correlate $\gamma_{\times}$ with $\phi$, the minor-axis reflection couples $x_{c}$ and $v_{c}$, and the rotation operation leaves all parameters independent. The bottom row shows results for a model with nonzero parameters ($\gamma_{\times}=0.01$, $\phi=5^{\circ}$, $x_{c}=y_{c}=1$, $v_{c}=15\ \mathrm{km\,s^{-1}}$), which introduce additional cross-correlations, particularly for the minor-axis reflection.
On the second issue of creating an averaged model versus separately computing a $\chi^{2}$ for the individual fields produced by each reflection, we find that summing over the pixels of the fields separately underestimates the error due to the individual fields not being truly independent. Consider the case of the reflection across the major axis alone. The residual is the difference between data and model, but the model itself is built from the data. If we look at a simplified, two-pixel image, with one upper and one lower pixel, identified with subscripts $U$ and $L$ respectively, the $\chi^{2}$ value (using unit uncertainty in the data) would be
@@
Alternatively, rather than reflecting the data, the operation can be thought of as a folding of the data across its axis, demonstrating that the number of true comparisons is only half the pixels. The same concept holds true for the reflection across the minor axis. By summing over the entirety of pixels in each transformed field separately, the number of independent data points is artificially inflated. Instead, we choose to average over two transformations, which creates a unique comparison for each pixel in the image.
Creating an averaged model produces the exact same result as computing separate $\chi^{2}$ values and manually accounting for the double-counting of pixels. We thus adopt the averaged model for simplicity while preserving the full information content of both symmetry operations.
III.5.2 Photometric Fitting
@@
where $M_{ij}=\sum_{x,y}x^{i}y^{j}I(x,y)$ denotes the raw moment, and $I(x,y)$ is the intensity at pixel coordinates $(x,y)$.
@@
With the axis ratio, we can calculate a model value for the photometric position angle from Equation 18. The axis ratio measurement introduces uncertainty in the model value, which must be accounted for. The uncertainty in the model term is calculated through standard propagation of errors
@@
Combining this term in quadrature with the uncertainty in the observed position angle gives the total error in the photometric angle. The uncertainties in observed position angle and axis ratio are determined directly from the distribution of position angles and axis ratios in our data sample, as shown in Section IV.1. This uncertainty effectively weights the extent to which the position angle contributes to the overall fit. It was shown in DiG21 that as this uncertainty increases, the scatter in the fitted shear trends to the results from a kinematic-only fit. A $\chi^{2}$ value is then calculated
@@
III.6 Data Selection
III.6.1 Idealized Mock Data
To characterize the performance of the MIRoRS method, we first create a set of idealized mock data for both the velocity and intensity fields. The intensity field is described by an exponentially decaying profile,
@@ III.6.1 Idealized Mock Data
For the velocity field, we model an infinitely thin disk with an arctangent profile with the major axis aligned along the $x$-axis,
@@
where $r_{0}$ is the scale radius independent of the intensity, $V_{\text{max}}$ is the asymptotic velocity, and $V_{c}$ is the central velocity of the galaxy.
@@
Both fields are then interpolated from the source-plane coordinates onto a grid that has been transformed by a shear parameter, $\gamma_{\times}$, major-axis position angle, $\phi$, and inclination $i$.
III.6.2 Illustris TNG
After characterization with idealized data, we then analyze a set of 358 subhalos from a $z=0$ snapshot of the Illustris TNG100-1 (Nelson et al., 2019) magnetohydrodynamical simulations. These halos were originally analyzed with parametric methods by DW23, and we refer the reader to their methods for an in-depth discussion on data selection. The general selection criteria required all samples to have a total mass within 10% of $1.4\text{\times}{10}^{12}$M_{\odot}$, with $\geq 40\%$ ($\leq 30\%$) of the stellar mass to be within the disk (bulge).
The subhalos were all oriented according to the principal axis of the moment of inertia tensor calculated from the stellar particles. The longest principal axis was aligned along the coordinate $x$-axis as the major axis, while the second-largest principal axis was aligned as the minor axis. The galaxies were also arranged to have a clockwise rotation.
III.7 Outlier Rejection
For a well-fit model, residuals are expected to be primarily due to uncorrelated noise. Spatially correlated patterns in residuals can be the result of a poorly fit parameter or from unmodeled physical effects such as spiral arms, bars, or warps. We use a Global Moran’s $I$ test for spatial autocorrelation (Moran, 1948) to quantify the extent to which these patterns are present in the residuals, identifying halos with a poor fit. Examining residuals for spatial autocorrelation is not a new technique (Goodchild, 1986), but has been largely limited to fields outside of astronomy. In using the Moran’s $I$ test, we seek to quantify the effectiveness of our modeling, but also to explore the utility of spatial autocorrelation in characterizing residuals.
To implement the Moran’s $I$ test, we first compute the residuals by subtracting the model values from the observed values. These residuals are then normalized by their combined errors, resulting in a standardized residuals image. The Moran’s $I$ statistic is then calculated to measure spatial autocorrelation, using the package PySAL.esda (Rey and Anselin, 2007). Specifically, we use the Queen contiguity method (named for the directions a Queen may move in chess) to create a spatial weight matrix, where each element $w_{ij}$ is 1 if locations $i$ and $j$ are neighbors (sharing a border or vertex) and 0 otherwise. The Moran’s $I$ statistic is given by